OpenStudy (anonymous):

Can there be three angles found for cosѲ= - 3/7 if 0°

OpenStudy (ikram002p):

whats the condition ? if 0° <or = 0 <or= 360 can u make it clear ?

OpenStudy (anonymous):

if the degree is more or equal than zero but less or equal to 360

OpenStudy (ikram002p):

u mean the angle should be with the range \(\large 0 \leq \theta \leq 2\pi\)

OpenStudy (anonymous):

0≤θ≤360 yea

ganeshie8 (ganeshie8):

familiar with `unit circle` ?

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):